A charge, q =-5pC moves through a uniform B-field, B = 0.35 T, with a velocity, v 200 m/s, at right angles to the field. Determine: (a) Fm on the charge; (b) mass of the charge if a 5x 10 m/s.

please I need help
Answer #1

Given q=-5mu C=-5*10^{-6}C

magnetic field, B 0.351.

velocity, v-200m/s

angle between magnetic field and velocity ,heta =90^{circ}

------------

Use Formula F=qvBsin heta

F--5 10200m/s 0.35T sin90

F=-5 * 10-6 * 200 * 0.35 * 1

(a)ANSWER: F =-3.5 * 10-4N

negative sign just shows that charge will move in opposite direction as in the case of positive charge

=======================

Use Formula F=ma

3.5 * 10-4 = m *-5 * 104m/s2

3.5 104-m*5 104

3.5 * 10-4/6 * 104) 772

(b)ANSWER: m=7*10^{-9}kg

==========================